We took an OLG measurement of the green PDH loop. It seems consistent with past measurements. I've added a trace for the the post-mixer lowpass, to show its contribution to the phase loss. (EDIT: updated with measured LPF TF)
I used this measured OLG and the datasheet laser PZT conversion factor to calibrate the control signal monitor into the AUX laser frequency noise, it looks consistent with the frequency noise measured via the PSL PLL (300 Hz/rtHz @ 100Hz). Above a few tens of kHz, the control signal measurement is all analyzer noise floor, due to the fourth order 70kHz lowpass after the mixer (the peaks change height significantly depending on the analyzer input range, so I don't think they're on the laser). Gautam will follow up with more detailed measurements of both the error and control signals as he noisebudgets, this was just intended as a quick consistency check.